Please help me with this problem I am having a diffcult time with it, Than you so much.
The population P in 1999 for a state given along with r, its annual percentage rate of continuous growth
P= 21 millions, r= 1.9%
(a) Write the formula f(x) = P e^yx, where r is in the decimal notation, that models the population in millions x years after 1999.
(b) Estimate the population in 2020
F(x) = ?
The population in 2020 will be approximately how many million?
(round to the nearest tenth as needed)
Any help is greatly appreciated!!!

Answer by ankor@dixie-net.com(22740)

The population P in 1999 for a state given along with r, its annual percentage rate of continuous growth
P = 21 millions, r= 1.9%
(a) Write the formula f(x) = P e^rx, where r is in the decimal notation, that models the population in millions x years after 1999.
f(x) = 21*e^.019x
:
(b) Estimate the population in 2020, (that's 21 yrs after 1999)
F(x) = 21*e^(.019*21)
f(x) = 21*e^.399
Find e^.399 on a calc
f(x) = 21 *1.4903
f(x) = 31.3 million in 2020